OpenStudy (anonymous):

Simplify; radical 18x + radical 2x = ?

OpenStudy (anonymous):

\[\sqrt{18x} + \sqrt{2x}\], and how would you start, any ideas?

OpenStudy (anonymous):

Simplify what's inside?

OpenStudy (anonymous):

yes so you can rewrite the 18 x radical as \[\sqrt{9 * 2x}\] and then the sqrt of 9 will come out the front. and then you would have = >?